A car moving at 150km/hr is approaching a stationary police car whose radar speed detector operates at a frequency of 15 Ghz.what frequency change is found by the speed detector?

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Vaidehi
 
Vr=150km/h * 5/18 
Vr = 41.66m/s. 

F = ((V-Vr)/(V+Vs))*Fo. 
F=((343-41.66) / (343+0))*15  = 13.178 Hz 

Change = 15 - 13.178  = 1.822 GHz.
 
Regards
Arun(askIITians forum expert)